SEARCHES FOR β-ALANINE AND PYRIMIDINE TOWARD ORION-KL

\underline{R. FUJIMORI}, K. KAWAGUCHI, Faculty of Science, Okayama University, Okayama 700-8530, Japan; T. NAKAJIMA, H. OGAWA, Department of Physical Science, Graduate School of Science, Osaka Prefecture University,Osaka 599-8531, Japan.

Watanabe et al.^{a} have searched for α-alanine in Ori-KL with the Nobeyama 45-m radio telescope. Blagojevic et al.^{b} expected that β-alanine may have a larger abundance than α-alanine in interstellar space. In meteorites the abundance ratio of glycine, β-alanine, α-alanine is reported to be 1 : 2.35 : 0.26. Pyrimidine may be related to the nucleic acid base such as uracil and searched by Kuan et al.^{c} In the present study, we carried out deep searches for β-alanine and pyrimidine toward Ori-KL by using a new waveguide-type sideband-separating receiver in the 3 mm band installed at Nobeyama 45-m telescope. Low noise characteristic of the receiver made it possible to high sensitive observation in wide freqency range. The rotational transitions of β-alanine and pyrimidine were not detected with upper limit abundances of 2.9×10^{14}cm^{−2}, 8.8×10^{12}cm^{−2}, respectivity. The upper limit for pyrimidine is an order of magnitude lower than the previous^{c}. Many un-identified lines have been detected and the carriers will be discussed.

aS. Watanebe, S. Kubota, K. Kawaguchi, Y. Kasai, and T. Momose, 61^{st} Ohio State University International Symposium on Molecular Spectroscopy, WH11 (2006).

bV. B. Blagojevic, S. Petrie, and D. K. Bohme, Mon. Not. R. Astron. Soc. 339, L7 (2003).

cY. J. Kuan, C. H. Yan, S. B. Charnley, Z. Kisiel, P. Ehrenfreund, and H. C. Huang, Mon. Not. R. Astron. Soc. 345, 650 (2003).